### v1

# Comparison of Structure-Preserving Methods for the Cahn-Hilliard-Navier-Stokes Equations

11institutetext: Jimmy Kornelije Gunnarsson 22institutetext: Centre for Mathematical Sciences, Lund University, Box 117, 22100 Lund, Sweden 22email: jimmy_kornelije.gunnarsson@math.lu.se 33institutetext: Robert Klöfkorn 44institutetext: Centre for Mathematical Sciences, Lund University, Box 117, 22100 Lund, Sweden 44email: robertk@math.lu.se
Comparison of Structure-Preserving Methods for the Cahn-Hilliard-Navier-Stokes Equations
[…]
1 Introduction
Consider the spatial domain $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, with Lipschitz continuous boundary $\partial\Omega$ and the time interval $(0,T]$ for $T\in\mathbb{R}^{+}$. The Cahn-Hilliard (CH) equations model phase separation in binary mixtures for a phase-field $\psi:\Omega\times(0,T]\to[-1,1]$ following:
$\partial_{t}\psi+\nabla\cdot(\psi\mathbf{u})-\nabla\cdot Pe^{-1}(M(\psi)\nabla\upsilon)$ $=$ $0,$ (1)
$\upsilon+Cn^{2}\Delta\psi-W^{\prime}(\psi)$ $=$ $0,$ (2)
where $Pe>0$ is the Peclet number, $\boldsymbol{u}$ is an advection field, $\upsilon$ is the chemical potential, $W(\psi):=\frac{1}{4}(1-\psi^{2})^{2}$ is the double-well potential, and $M(\psi):=\max{\{1-\psi^{2},0\}}$ is the mobility function. We equip Eq.(1) with homogeneous Neumann boundary conditions $\boldsymbol{n}\cdot\nabla\psi=0$ and $\boldsymbol{n}\cdot\nabla\upsilon=0$ on $\partial\Omega$.
We present below a special case of a fundamental theorem for boundedness.
Theorem 1.1(Boundedness Elliott and Garcke (1996))
Suppose that $\Omega$ is convex and that the initial phase-field satisfies $\psi^{0}\in H^{1}(\Omega)$ and $\psi^{0}||_{L^{\infty}(\Omega)}\leq 1$. Then, if the mobility function $M(\psi)$ is defined as above, and $\int_{\Omega}M(\psi^{0})+W(\psi^{0})<C$ for $C>0$, then the weak solution $\psi\in H^{1}(\Omega)$ of the CH equation satisfies a weak maximum principle, i.e., $\psi||_{L^{\infty}(\Omega)}\leq 1$ for $t\in(0,T]$.
While Theorem 1.1, proven in (Elliott and Garcke, 1996, Thm. 1), guarantees a weak maximum principle at the continuum level for a degenerate mobility, achieving discrete boundedness remains a significant challenge. The standard Finite Element Method (FEM), symmetric weighted interior penalty (SWIP), and symmetric interior penalty Galerkin (SIPG) discretizations fail to preserve boundedness without additional stabilization Gunnarsson and Klöfkorn (2026), even when using the degenerate mobility $M$. Recent work, however, has demonstrated that boundedness can be numerically achieved through carefully designed limited Galerkin schemes, including FEM-L for FEM, and for DG SIPG-L, and SWIP-L Gunnarsson and Klöfkorn (2026) with degenerate CH.
In this work, we extend these structure-preserving methods by introducing new mobility fluxes using the Discontinuous Galerkin (DG) formulations with intersection-wise mobility treatments. These extend the SWIP-L and SIPG-L methods introduced in Gunnarsson and Klöfkorn (2026), by allowing for better control of the coercivity-stability balance.
Boundedness of the CH equations is in particular important when coupled to the incompressible Navier-Stokes (NS) equations:
$\rho(\psi)\big(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}\big)+\boldsymbol{J}\cdot\nabla\boldsymbol{u}+\nabla\cdot\big(P\mathbb{I}-\mu(\psi)D(\boldsymbol{u})\big)$ $=$ $\frac{1}{WeCn}\upsilon\nabla\psi$ (3)
$\nabla\cdot\mathbf{u}$ $=$ $0$ (4)
where $\rho(\psi):=\frac{1+\psi}{2}\rho_{1}+\frac{1-\psi}{2}\rho_{2}$ is the density, $\boldsymbol{J}:=\frac{\rho_{2}-\rho_{1}}{2}Pe^{-1}M(\psi)\nabla\upsilon$ is the mass flux, $P$ is the pressure, $\mu(\psi):=\frac{1+\psi}{2}\mu_{1}+\frac{1-\psi}{2}\mu_{2}$ is the viscosity, and $D(\boldsymbol{u})=\frac{1}{2}(\nabla\boldsymbol{u}+\nabla\boldsymbol{u}^{T})$ is the rate of deformation tensor, where $\mathbb{I}$ denotes the identity tensor. Moreover, $\rho_{j}$ and $\mu_{j}$ are the characteristic densities and viscosities for phases $j=1,2$, respectively. We equip the boundary conditions $\mathbf{n}\cdot\boldsymbol{u}|_{\partial\Omega}=\mathbf{n}\cdot\nabla P|_{\partial\Omega}=0$. We note that the expected bounds for the physical mass density $\rho\in[\rho_{1},\rho_{2}]$ follows directly from boundedness of the phase-field $\psi\in[-1,1]$.
The main contributions of this paper are: (i) the development of SWIPD-L and SIPGD-L methods with a parametrized mobility flux for enhanced stability, (ii) a proof of coercivity for the generalized trilinear form for SWIPD and SIPGD schemes, and (iii) numerical validation demonstrating optimal convergence rates and structure preservation for both standalone CH and coupled CHNS systems. We compare the performance of SIPG-L, SWIP-L, SIPGD-L, and SWIPD-L methods through extensive numerical experiments, including $hp$-adaptive formulations for our new schemes.
2 Discretization
In this section, we consider the spatial and temporal discretization of the CHNS equation system. Our main focus is to motivate the design of the SWIPD-L and SIPGD-L methods and to provide a proof of coercivity for the resulting trilinear form, while the remaining details follow a similar procedure as in Gunnarsson and Klöfkorn (2026).
Let the spatial domain $\Omega$ be partitioned into a union of $M$ non-intersecting elements $K$ forming a mesh ${\mathcal{T}_{h}}=\cup_{i=1}^{M}K_{i}$. Then we denote by $\Gamma_{i}$ with unit normal $\boldsymbol{n}$ the set of all intersections between two elements of the grid $\mathcal{T}_{h}$}, and the set of all intersections, also with the boundary of the domain $\Omega$, is denoted by $\Gamma$. For an intersection $e\in\Gamma$ we denote the adjacent elements with $K^{-}_{e}$ and $K^{+}_{e}$ ($K^{-}_{e}=K^{+}_{e}$ for $e\in\Gamma\cap\Gamma_{i}$) and define
$h_{e}:=\frac{2|K^{-}_{e}||K^{+}_{e}|}{|e|(|K^{-}_{e}|+|K^{+}_{e}|)}\quad\forall e\in\Gamma.$ (5)
The global mesh width is then defined as $h=\max_{e\in\Gamma}h_{e}$. For the time discretization we consider a uniform partition of the time interval $[0,T]$ with time increment $\tau=\frac{T}{N}$ for some $N\in\mathbb{N}$ and denote the time levels with $t_{n}=n\cdot\tau$ for $n=0,1,\ldots,N$. For brevity we will also denote time-dependent variables with a subscript $n$, i.e., $\psi^{n}=\psi(\cdot,t_{n})$.
Following standard FEM notation we consider a general order FEM formulation for the function space of trial and test functions:
$V_{h}^{p}=\{\varphi\in L^{2}({\mathcal{T}_{h}}):\varphi|_{K}\in\mathbb{P}^{p}(K),\forall K\in{\mathcal{T}_{h}}\},$ (6)
where $\mathbb{P}^{p}({\mathcal{T}_{h}})$ denotes a polynomial space of order at most $p$ on the grid $\mathcal{T}_{h}$}. We also permit $p$-adaptivity over the mesh $\mathcal{T}_{h}$}, and as such, the polynomial order $p$ may vary across elements $K$, and we denote the local polynomial order as $p_{K}$. We denote the maximal permitted polynomial order as $p_{\max}$, and the minimal order as $p_{\min}$. We describe the application further in Sec. 2.2.
Before proceeding, we introduce operators ${\cdot}$$\{{\cdot}\}$ and $\langle{\cdot}\rangle$ for $e\in\Gamma_{i}$ as
[…]
for some $\varphi$, where $\{{\cdot}\}$ and $\langle{\cdot}\rangle$ denote the arithmetic and harmonic averages, respectively. To simplify the notation $\varphi^{\pm}:=\varphi_{|K_{e}^{\pm}}$ is also used frequently. Moreover, we introduce the subscript $\varphi_{\oplus}:=\max{\{0,\varphi\}}$ and $\varphi_{\ominus}:=\min{\{0,\varphi\}}$ to denote the positive and negative restriction of a function, respectively. This notation will in particular be utilized for upwinding.
2.1 Discontinuous Galerkin
Given a degenerate mobility function $M:V_{h}^{p}\to[0,1]$, we introduce the trilinear form $b:L^{\infty}({\mathcal{T}_{h}})\times V_{h}^{p}\times V_{h}^{p}\to\mathbb{R}$:
$b(M(\psi_{h}),\upsilon_{h},\varphi)$ $=\int_{\mathcal{T}_{h}}M(\psi_{h})\,\nabla\upsilon_{h}\cdot\nabla\varphi\,dx+\sum_{e\in\Gamma}\int_{e}\frac{\eta_{e}\Lambda_{e}(M(\psi_{h}))}{h_{e}}[{\upsilon_{h}}][{\varphi}]$ (7)
$-F(M(\psi_{h}),\upsilon_{h})\,[{\varphi}]-F(M(\psi_{h}),\varphi)\,[{\upsilon_{h}}]ds,$
where $\eta_{e}\geq 0$ is a penalty parameter over $e$, $F$ is the mobility flux, and $\Lambda_{e}(M(\psi_{h}))$ is a mobility flux function. We examine the coercivity properties arising from the degeneracy of $M$. We consider the following formulations for $F$:
$F(M(\psi_{h}),\upsilon_{h})=\langle{M(\psi_{h})}\rangle\{{\nabla\upsilon_{h}}\},\quad F(M(\psi_{h}),\upsilon_{h})=\{{M(\psi_{h})\nabla\upsilon_{h}}\},$ (8)
corresponding to a weighted average with a harmonic average for the mobility flux $\langle{M(\psi_{h})}\rangle$ and an arithmetic average $\{{M(\psi_{h})\nabla\upsilon_{h}}\}$, which correspond to the SWIP and SIPG discretization, respectively. We note that the SWIP formulation with $\langle{M(\psi_{h})}\rangle$ follows from choosing the weights $w^{\pm}=\frac{M(\psi_{h})^{\mp}}{M(\psi_{h})^{+}+M(\psi_{h})^{-}}$, as suggested in Ern et al. (2008) for treating a diffusion-type formulation for DG.
Lemma 1(Trace inequality, Rivière (2008))
Consider an intersection $e\in\Gamma_{i}$ shared by two elements $K^{-}_{e}$ and $K^{+}_{e}$. Then for $K\in{\mathcal{T}_{h}}$ there exists a constant $C_{K}>0$ independent of the mesh width $h$ such that for all $\varphi_{h}\in V^{p}_{h}$ for $p>0$ and $e\in\Gamma_{i}$ the following inequality holds:
[…]
where, in the special case of $K_{e}$ being a quadrilateral or triangle, $C_{K_{e}^{\pm}}=\frac{p_{K}(p_{K}+d-1)}{d}$ where $p_{K}$ is the local polynomial degree (see further treatments and formulations in Ainsworth and Rankin (2009) and Epshteyn and Rivière (2007)).
Lemma 2(Trace inequality for mixed polynomial order)
Under the same setting as in Lemma 1 and for $\varphi_{h}\in V_{h}^{p}$ the following inequality holds for all $e\in\Gamma_{i}$:
[…]
where for the special case when $K_{e}$ is a quadrilateral or triangle:
$C_{e}=\underset{{p\in\{p_{K^{+}_{e}},p_{K^{-}_{e}}\}}}{\max}\frac{p(p+d-1)}{d}.$ (11)
Proof
The proof follows from the triangle inequality and the trace inequality in Lemma 1:
[…]
then upon applying the trace inequality in Lemma 1 we obtain the desired result by then maximizing the quantity
$\frac{|e|}{|K_{e}^{+}|}C_{K_{e}^{+}}+\frac{|e|}{|K_{e}^{-}|}C_{K_{e}^{-}}\leq\frac{C_{e}}{h_{e}},$ (13)
for a joint inequality.
Theorem 2.1(Coercivity)
[…]
for some $C>0$ where $\upsilon_{h}||_{DG}^{2}\leq C\,b(M(\psi_{h}),\upsilon_{h},\upsilon_{h})$. Then the following coercivity condition holds:
$\eta_{e}\geq\max_{p\in\{K_{e}^{+},K_{e}^{-}\}}p(p+d-1),\quad\Lambda_{e}(M(\psi_{h}))\geq\tilde{M}_{e}^{2\alpha}(\psi_{h})\lambda^{\star},$ (14)
where for $\alpha\in[0,\frac{1}{2}]$ the maximum contrast $\lambda^{\star}$ is defined as
[…]
and the flux $\tilde{M}_{e}(\psi_{h})$ is defined as
$\tilde{M}_{e}(\psi_{h})=\langle{M(\psi_{h})}\rangle,\quad\tilde{M}(\psi_{h})=\max\{M(\psi_{h}^{+}),M(\psi_{h}^{-})\},$ (16)
for the SWIP and SIPG mobility fluxes, respectively.
Proof
The overall proof follows similarly to the one in (Gunnarsson and Klöfkorn, 2026, Theorem 3.2), with the main difference being the treatment of the mobility term $M(\psi_{h})$ for the first Cauchy-Schwarz inequality over the intersections $e\in\Gamma_{i}$. In particular we note that for $e\in\Gamma_{i}$ we have
[…]
Following a similar approach to the proof in Gunnarsson and Klöfkorn (2026), which considered $\alpha=0$, we generalize to a free variable $\alpha\in[0,\frac{1}{2}]$ to obtain the desired result by also applying the modified trace inequality in Lemma 2 and maximizing the resulting quantity for a joint inequality over the intersection $e$.
For the remainder of this study we will consider the regularization $M_{\delta}(\psi_{h})=\max\{M(\psi_{h}),\delta\}$ for some small $\delta>0$ to ensure the coercivity condition in Theorem 2.1 is satisfied, and we will simply denote this regularized mobility as $M(\psi_{h})$ for brevity and we pick $\delta=10^{-20}$ (practically 0) for the numerical simulations.
Remark 1(mobility flux)
It is complicated to estimate the maximum contrast $\lambda^{\star}$ for a sharp estimate. For simplicity, it is sufficient to consider the mobility flux as $\Lambda_{e}=\beta\tilde{M}_{e}^{2\alpha}(\psi_{h})$ for some constant a-priori estimate $\beta\geq\lambda^{\star}$. For instance, $\beta=5$ was considered in the numerical experiments in Gunnarsson and Klöfkorn (2026).
Corollary 1(Case $V_{h}^{0}$)
For the case of $V_{h}^{0}$ the coercivity condition in Thm. 2.1 reduces to
$\eta_{e}\geq 1,\quad\Lambda_{e}(M(\psi_{h}))\geq\tilde{M}(\psi_{h}).$ (18)
Proof
Since for $\varphi\in V_{h}^{0}$ then $\nabla\varphi=0$ and thus only the penalty term remains in the trilinear form $b$ in Eq.(7), and consequently the coercivity condition reduces to $\eta_{e}\Lambda_{e}(M(\psi_{h}))\geq\tilde{M}(\psi_{h})$, which is satisfied for the stated conditions.
Similarly to Eq.(7) we introduce the following bilinear form for the discretization of the Laplacian operator in Eq.(1):
$a(\psi_{h},\varphi)$ $=\int_{\mathcal{T}_{h}}\nabla\psi_{h}\cdot\nabla\varphi\,dx+\sum_{e\in\Gamma}\int_{e}\frac{\eta_{e}^{\Delta}}{h_{e}}[{\psi_{h}}][{\varphi}]$ (19)
$-\{{\nabla\psi_{h}\cdot\boldsymbol{n}}\}[{\varphi}]-\{{\nabla\varphi\cdot\boldsymbol{n}}\}[{\psi_{h}}]ds,$
where $\eta_{e}^{\Delta}\geq\eta_{e}$ is a local penalty parameter for the Laplacian operator. We introduce the superscript $\Delta$ to distinguish the penalty parameter for the Laplacian from the one for the mobility term in Eq.(7), and in particular, to lump the constants in practice so that $\eta_{e}^{\Delta}=\beta\eta_{e}$ for some $\beta\geq 1$ following Rmk. 1.
When we consider advection coupling we introduce the standard upwinding for the advection term:
$c(\boldsymbol{u},\psi_{h},\varphi)=\int_{\mathcal{T}_{h}}\boldsymbol{u}\cdot\nabla\varphi\,\psi_{h}\,dx-\sum_{e\in\Gamma_{i}}\int_{e}\left(\{{\boldsymbol{u}\cdot\mathbf{n}^{+}}\}_{\oplus}\psi_{h}^{+}+\{{\boldsymbol{u}\cdot\mathbf{n}^{+}}\}_{\ominus}\psi_{h}^{-}\right)[{\varphi}]ds,$ (20)
Thus, we have introduced all the DG-specific bilinear and trilinear forms for the spatial discretization of the CH equations. In the following, we consider the time discretization and coupling to the NS equations.
To treat the non-linear term $W$ we introduce the Eyre splitting:
$\Phi^{+}(\psi_{h})=\psi_{h}^{3},\quad\Phi^{-}(\psi_{h})=-\psi_{h},$ (21)
where clearly $W^{\prime}(\psi_{h})=\Phi^{+}(\psi_{h})+\Phi^{-}(\psi_{h})$. Below we present our proposed schemes for the CH equations:
$\frac{1}{\tau}\left\langle\psi_{h}^{n}-\psi_{h}^{n-1},\varphi\right\rangle-c(\tilde{\boldsymbol{u}}^{n-\frac{1}{2}},\psi_{h}^{n},\varphi)+Pe^{-1}b(M(\psi_{h}^{n}),\upsilon_{h}^{n},\varphi)$ $=$ $0,$ (22)
$\left\langle\upsilon_{h}^{n},\xi\right\rangle-\left\langle\Phi^{+}(\psi_{h}^{n})+\Phi^{-}(\psi_{h}^{n-1}),\xi\right\rangle-Cn^{2}a(\psi_{h}^{n},\xi)$ $=$ $0,$ (23)
for $\xi,\varphi\in V_{h}^{p}$ and $\tilde{\boldsymbol{u}}^{n-\frac{1}{2}}$ being the velocity field at time step $n-\frac{1}{2}$. We note that the mobility term, and penalty weights, are treated implicitly in time. The main difference between the proposed schemes is the mobility flux for the penalty term, where we consider both a harmonic average and an intersection maximum as described in Eq.(7). The outline and algorithm of the scheme are very similar to those detailed in Gunnarsson and Klöfkorn (2026), as it is simply an algebraic reformulation of the same scheme with a different mobility flux for the penalty term, therefore we do not repeat the details here.
For the Navier-Stokes coupling we follow exactly the same formulation as in Gunnarsson and Klöfkorn (2026) with a post-processing step. Our modification includes projecting the FEM velocity field $\boldsymbol{u}$ to a Brezzi-Douglas-Marini space of first order to obtain a divergence-free velocity field $\tilde{\boldsymbol{u}}$.
Definition 1(Phase-field mass)
We refer to the quantity
$m_{\psi_{h}}=\frac{1}{|{\mathcal{T}_{h}}|}\int_{\mathcal{T}_{h}}\psi_{h}\,dx$ (24)
with $|{\mathcal{T}_{h}}|=\int_{\mathcal{T}_{h}}1\,dx$ as the phase-field mass.
Definition 2(Discrete energy)
We refer to the quantity:
$\mathcal{E}[\psi_{h},\boldsymbol{u}]=\frac{1}{WeCn}\int_{\mathcal{T}_{h}}W(\psi_{h})+\frac{Cn^{2}}{2}a(\psi_{h},\psi_{h})\,dx+\int_{\mathcal{T}_{h}}\rho(\psi_{h})\frac{|\boldsymbol{u}|^{2}}{2}\,dx,$ (25)
as the discrete energy.
Following similar arguments as in Gunnarsson and Klöfkorn (2026) and references therein, the provided schemes are energy dissipative and mass conservative on the discrete level for the quantities defined in Definitions 1 and 2.
2.2 Limiter and $hp$-adaptivity
Figure 1: Example of a scaling limiter for local $\mathbb{P}^{2}$ phase-field $\psi_{h}$
A Zhang-Shu scaling limiter Zhang and Shu (2010) is applied to $\psi_{h}$, with quadrature evaluation for local polynomial order $p_{K}\geq 1$. The limiter structure is outlined in Fig. 1 and described in Gunnarsson and Klöfkorn (2026). Following the notation in Gunnarsson and Klöfkorn (2026) we denote by SWIPD-L and SIPGD-L the methods corresponding to the mobility flux with $\alpha=\frac{1}{2}$ in Theorem 2.1, while the methods with $\alpha=0$ are denoted by SWIP-L and SIPG-L, respectively. We note that the mobility flux with $\alpha=\frac{1}{2}$ corresponds to a more diffusive flux, which facilitates treatment over the intersections $e$, at the potential cost of a larger global contrast $\lambda^{\star}$.
For $h$-adaptivity we follow the formulation provided in Gunnarsson and Klöfkorn (2026) and we consider an indicator function $H(\psi_{h}):=\frac{(1-\psi^{2})}{4}$ and refinement for elements $K$ where $H(\psi_{h})<0.0525$, and coarsening is suggested when $H(\psi_{h})>0.15$. Meanwhile, for $p$-adaptivity, coarsening is performed for elements $K$ where $H(\psi_{h})>0.075$ and refinement follows similarly to $h$-adaptivity. Fig. 2 shows an example of $hp$-adaptivity. We note that the thresholds for refinement and coarsening are somewhat arbitrary and can be tuned for better performance.
Figure 2: Ex. 3 with hphp-adaptivity for $p_{\min}=1$ and $p_{\max}=2$ at time $t=0$.
In particular, we aim to isolate elements $K$ with low-order polynomials that satisfy $\psi_{h}||_{L^{\infty}(K)}\lessapprox 1$ to ensure that coercivity is not broken by populating elements with low-order polynomials with values close to the bounds, which can ideally lead to a smaller contrast $\lambda^{\star}$ and less penalization in the scheme. Such effects are examined in Section 3, where we compare the performance of $h$-adaptivity and $hp$-adaptivity for the proposed schemes. Moreover, to allow $p_{\min}=0$ we consider $\eta_{e}=\underset{p\in\{p_{K^{+}_{e}},p_{K^{-}_{e}}\}}{\max}\max\left\{p(p+d-1),1\right\}$.
3 Numerics
In this section, we present numerical results for the proposed schemes for both the standalone CH equations and the coupled CH-NS system. We first consider Ex. 1 to verify the optimal convergence rates in 2D and 3D. We then investigate the performance of the proposed schemes for merging droplets in Ex. 2, and finally consider the rotating droplet configuration in Ex. 3 for the coupled CH-NS system to demonstrate the structure-preserving properties and $hp$-adaptivity. The numerical simulations are implemented in Dune-Fem Bastian et al. (2021); Dedner and Klöfkorn (2021) at version $2.11$ with its Python interface Dedner et al. (2020).
Example 1(Artificial trigonometric solution)
Consider the domain $\Omega=[0,1]^{d}$ and the following artificial solution to the CH equations $\psi(\boldsymbol{x},0)=A\prod_{i=1}^{d}\cos(4\pi\boldsymbol{x}_{i})$ where $A\in(0,1]$ is an amplitude parameter and $\boldsymbol{x}_{i}$ is the $i$-th component of the spatial coordinate $\boldsymbol{x}$. We then consider the following forcing term $S(\boldsymbol{x},t)=Pe^{-1}\nabla\cdot(M(\psi)(Cn^{2}\Delta\psi-W^{\prime}(\psi)))$ added to the right-hand side of Eq.(22), where $Cn=0.1$, $Pe=0.3$ are the Cahn and Peclet numbers respectively. The simulation is run for $t\leq T=0.01$.
Table 1: Ex. 1: $L^{2}$ and $H^{1}$ errors and EOC for $A=0.1$ and $A=0.6$ and $\tau=10^{-(2+p)}\cdot 2^{-\frac{N-32}{32}}$ for $p\in\{1,2\}$ and $N$ being the number of elements in each axis.
$A=0.1$ $A=0.6$
$L^{2}$ Error $H^{1}$ Error $L^{2}$ Error $H^{1}$ Error
Scheme, $p$ NN Error EOC Error EOC Error EOC Error EOC
32 1.29e-03 1.01e-01 7.14e-03 6.05e-01
64 3.41e-04 1.92 5.04e-02 1.00 1.88e-03 1.92 3.02e-01 1.00
SIPG-L, 11 128 8.75e-05 1.96 2.52e-02 1.00 4.82e-04 1.97 1.51e-01 1.00
256 2.21e-05 1.98 1.26e-02 1.00 1.22e-04 1.98 7.56e-02 1.00
32 2.19e-05 5.10e-03 1.32e-04 3.06e-02
64 2.72e-06 3.01 1.28e-03 2.00 1.63e-05 3.01 7.66e-03 2.00
SIPG-L, 22 128 3.40e-07 3.00 3.19e-04 2.00 2.04e-06 3.00 1.91e-03 2.00
256 4.25e-08 3.00 7.98e-05 2.00 2.55e-07 3.00 4.79e-04 2.00
32 1.29e-03 1.01e-01 7.14e-03 6.05e-01
64 3.41e-04 1.92 5.04e-02 1.00 1.88e-03 1.92 3.02e-01 1.00
SWIP-L, 11 128 8.75e-05 1.96 2.52e-02 1.00 4.82e-04 1.97 1.51e-01 1.00
256 2.21e-05 1.98 1.26e-02 1.00 1.22e-04 1.98 7.56e-02 1.00
32 2.19e-05 5.10e-03 1.32e-04 3.06e-02
64 2.72e-06 3.01 1.28e-03 2.00 1.63e-05 3.01 7.66e-03 2.00
SWIP-L, 22 128 3.40e-07 3.00 3.19e-04 2.00 2.04e-06 3.00 1.91e-03 2.00
256 4.25e-08 3.00 7.98e-05 2.00 2.55e-07 3.00 4.79e-04 2.00
32 1.29e-03 1.01e-01 7.14e-03 6.05e-01
64 3.41e-04 1.92 5.04e-02 1.00 1.88e-03 1.92 3.02e-01 1.00
SIPGD-L, 11 128 8.75e-05 1.96 2.52e-02 1.00 4.82e-04 1.97 1.51e-01 1.00
256 2.21e-05 1.98 1.26e-02 1.00 1.22e-04 1.98 7.56e-02 1.00
32 2.19e-05 5.10e-03 1.32e-04 3.06e-02
64 2.72e-06 3.01 1.28e-03 2.00 1.63e-05 3.01 7.66e-03 2.00
SIPGD-L, 22 128 3.40e-07 3.00 3.19e-04 2.00 2.04e-06 3.00 1.91e-03 2.00
256 4.25e-08 3.00 7.98e-05 2.00 2.55e-07 3.00 4.79e-04 2.00
32 1.29e-03 1.01e-01 7.14e-03 6.05e-01
64 3.41e-04 1.92 5.04e-02 1.00 1.88e-03 1.92 3.02e-01 1.00
SWIPD-L, 11 128 8.75e-05 1.96 2.52e-02 1.00 4.82e-04 1.97 1.51e-01 1.00
256 2.21e-05 1.98 1.26e-02 1.00 1.22e-04 1.98 7.56e-02 1.00
32 2.19e-05 5.10e-03 1.32e-04 3.06e-02
64 2.72e-06 3.01 1.28e-03 2.00 1.63e-05 3.01 7.66e-03 2.00
SWIPD-L, 22 128 3.40e-07 3.00 3.19e-04 2.00 2.04e-06 3.00 1.91e-03 2.00
256 4.25e-08 3.00 7.98e-05 2.00 2.55e-07 3.00 4.79e-04 2.00
Table 2: Ex. 1 in 3D: EOC for $L^{2}$ and $H^{1}$ errors for $A=0.1$ and $\tau=10^{-3}\cdot 2^{-\frac{N-8}{8}}$.
$L^{2}$ $H^{1}$
Scheme $p$ NN Error EOC Error EOC
8 5.50e-03 3.48e-01
16 1.40e-03 1.97 1.76e-01 0.98
SIPGD-L 11 32 3.51e-04 1.99 8.76e-02 1.01
64 8.80e-05 2.00 4.37e-02 1.00
8 5.50e-03 3.48e-01
16 1.40e-03 1.97 1.76e-01 0.98
SWIPD-L 11 32 3.51e-04 1.99 8.76e-02 1.01
64 8.80e-05 2.00 4.37e-02 1.00
In Table 1, we present the $L^{2}$ and $H^{1}$ errors and EOC for Ex. 1 for the different schemes with amplitudes $A=0.1$ and $A=0.6$ and $\beta=5$. The results are consistent with the theoretically expected convergence rates. Additionally, SWIPD-L/SIPGD-L and SWIP-L/SIPG-L yield identical errors and EOC, which is consistent with the fact that the mobility flux does not affect the consistency of the scheme, but only the stability and coercivity. Similar behavior for SIPG-L and SWIP-L was also observed in Gunnarsson and Klöfkorn (2026). We also tested if the same convergence rates hold in 3D, and the results are presented in Table 2 for $A=0.1$ and $\tau=10^{-3}\cdot 2^{-\frac{N}{16}}$. The observed rates confirm the theoretical predictions.
For the remainder of the numerical experiments, we employ the SWIPD-L scheme with $\alpha=\frac{1}{2}$, as it yields a more diffusive flux and consequently improved consistency. Furthermore, we use $\beta=3$ for the mobility flux as in Rmk. 1.
Example 2(Merging droplets)
We consider the CH equations without advection, i.e., $\mathbf{u}(\cdot,\cdot)=\boldsymbol{0}$, in the domain $\Omega=[0,1]^{2}$. The initial condition is given by a smooth profile
[…]
where $r=0.2$ is the droplet radius, with central points $\mathbf{c}_{1}=(0.3,0.5)^{T}$ and $\mathbf{c}_{2}=(0.7,0.5)^{T}$, Cahn number $Cn=4h_{\min}$, and Peclet number $Pe^{-1}=3Cn$. The simulation is run for $t\leq T=0.4$.
Example 3(Rotating Merging Bubbles)
We consider the CHNS equations in the domain $\Omega=[-0.5,0.5]^{2}$, with a divergence-free initial velocity field
[…]
where $\boldsymbol{x}=(x_{1},x_{2})^{T}$, and $\chi=100$ is a scaling factor. The initial phase-field profile is
[…]
where $r_{1}=0.25$ and $r_{2}=0.15$ are the radii of the respective droplets, with central points $\mathbf{c}_{1}=(0.1,0.1)^{T}$ and $\mathbf{c}_{2}=(-0.15,-0.15)^{T}$, respectively. The following non-dimensional numbers are considered: Reynolds number $Re=1$, Cahn number $Cn=4h_{\min}$, Weber number $We=Cn^{-1}$, viscosities $\mu_{1}=\mu_{2}=1$, densities $\rho_{1}=100$ and $\rho_{2}=1$, and Peclet number $Pe^{-1}=3Cn$. The simulation is run for $t\leq T=0.2$.
For Exs. 2 and 3, we consider an $hp$-adaptive formulation with $h_{\min}=\frac{1}{256}$ and $h_{\max}=\frac{1}{128}$ using the SWIPD-L scheme and a prescribed maximum polynomial degree of $p_{max}$ and $p_{min}\geq 0$. For the time increments, following Gunnarsson and Klöfkorn (2026), we set $\tau=10^{-3}$ for Ex. 2 and $\tau=5\cdot 10^{-4}$ for Ex. 3. We simulate with different order levels of $p$ for $p$-adaptivity and also compare to only the $h$-adaptive formulation, i.e., $p=p_{max}$ for all elements $K\in{\mathcal{T}_{h}}$. We observe that convergence was not achieved for Ex. 2 with $p=2$ for the $h$-adaptive formulation; convergence was only obtained, in that case, when $\psi_{h}||_{L^{\infty}(K)}\lessapprox 1$, as was also stressed in Gunnarsson and Klöfkorn (2026) regarding careful selection of the initial data. To the best of the authors’ knowledge there is no clear proof or indication on how to find a safe initial value.
Figure 3: Ex. 2: Energy, mass, and boundedness for SWIPD-L.
Figure 3 shows the energy, mass, and boundedness for Ex. 2 using the $hp$-adaptive formulation of SWIPD-L with for various $p$-levels for $hp$-adaptivity. The energy is monotonically non-increasing, mass is conserved, and the phase-field remains bounded, consistent with the theoretical properties. The results are similar to those obtained using only $h$-adaptivity with $p=1$, and for $p_{\max}=1$ with $hp$-adaptivity, which suggests that the $hp$-adaptive formulation preserves accuracy alongside the $h$-adaptive formulation, while providing a significant reduction in complexity and degrees of freedom.
Figure 4: Ex. 3: Energy, mass, and boundedness for SWIPD-L.
Figure 4 shows that the energy is non-increasing, mass is conserved, and the phase-field remains bounded for Ex. 3 using the $hp$-adaptive formulation of SWIPD-L with and without $p$-adaptivity for the studied levels. The results are similar to those obtained using only $h$-adaptivity for the studied $p$-levels, which further indicates that the $hp$-adaptive formulation preserves physical accuracy alongside to the $h$-adaptive formulation, while providing a significant reduction in complexity and degrees of freedom, similarly to our discussion for Ex. 2. We also note that for $p=2$ the mass difference in Fig. 4 is slowly increasing due to the nonlinear tolerance $\epsilon$ (for these simulations we picked $\epsilon=5\cdot 10^{-16}$) having to be picked slightly larger to avoid solver stagnation. These minor mass drifts are a known issue with potential deviations scaling as the non-linear tolerance as noted in Gunnarsson and Klöfkorn (2026). Regardless, as is seen in Figs. 3 and 4 showing that the mass change is negligible in comparasion to the lines at $\Delta m=\pm\frac{T}{\tau}\epsilon$ for $\epsilon=2\cdot 10^{-16}$ for Ex. 2 and $\epsilon=5\cdot 10^{-16}$ for Ex. 3.
4 Summary and Outlook
We have proposed a new class of DG schemes for the CH equations with degenerate mobility has been proposed, which are provably coercive and stable under a certain condition on the mobility flux for the penalty term. The introduced mobility flux formulations for the penalty term include SIPGD-L, which employs the intersection maximum of the mobility, and SWIPD-L, which uses a harmonic average. These mobility fluxes yield similar estimates for the coercivity condition and accuracy as previous schemes. Numerical experiments illustrate the performance of the proposed schemes, demonstrating that they achieve optimal convergence rates and that the harmonic mobility flux provides potentially significant improvements in stability and accuracy compared to the arithmetic mean. An $hp$-adaptive formulation has also been presented, showing significant reductions in complexity and degrees of freedom compared to the $h$-adaptive formulation while maintaining similar accuracy and physical fidelity.
Future work includes extending the theoretical analysis to provide rigorous proofs of boundedness for the SWIPD-L scheme in the broken Sobolev (DG) setting, exploiting the connection between the harmonic average and the minimum to extend Theorem 1.1. Additionally, more challenging test cases, such as the rising bubble benchmark, will further validate the performance of the proposed schemes for coupled CHNS systems and assess adaptivity effects under more demanding conditions.